i need to solve for X: x/a+x/b=1 it's x over a plus x over b = 1, someone just provided the answer and that does not help. i need to see how you actually arrived at the answer.
Answer by lynnlo(4176)

OK,THIS ONE IS SIMPLE
PROBLEM:
X/A+X/B
X/A+X/B=0
1/A-1/B=0
X=AB/A+B